Does training in study skills make a difference for student success? 8 students received training (M1= 85, SS1= 100) and 8 students did not (M2= 80, SS2= 124). Test using α = .05  

Which is the null hypothesis, H0?

µ1 = 0

µ1- µ2 ≠ 0

µ1- µ2 = 0

µ1 ≠ 0

The degrees of freedom (total) (df) =

18

14

16

10

With alpha = .05, (two tail, nondirectional test), the critical t score boundaries =

+/- 2.145

+/- 2.179

+/- 2.120

+/- 2.776

s2p= (pooled variance) =

20

16

18

12

s(M1-M2) =

8

2

4

1

t = (test statistic) =

+1.75

+3.00

+1.25

+2.50

What is the correct result based on the data?

Reject Ho, result is Significant

Reject Ho, result is NOT Significant

Retain Ho, result is NOT Significant

Retain Ho,result is Significant

What type of error could you be making?

Type I

Type II

Does training in study skills make a difference?

No

Yes

The correct APA format for the result would include

p > .05

p < .05

I know this is a lot but it's all one problem. You dont have to answer all if you cannot, but it would be so much appreciated.
Answer #1

ni -8 XI - Given : 85 SS 100 n2-8 X2 = 80 SS2 = 124 Ho: Mi-42=0 Hi: li-42 € 0 df = nitn2-2 = 8 + 8-2 14 critical value at x =since, iti > tcritical, Wale critical, we reject Ho and conclude that result is significant type of Error = Type - training i